OpenStudy (anonymous):

factor x^2-2x-3 whats the answer

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what two numbers multiply to be 3 and add to be 2?

OpenStudy (amistre64):

i take it you are done with the other question we were working on?

OpenStudy (anonymous):

@amistre. what is the greatest common factor of 275?

OpenStudy (amistre64):

id say 275 :)

OpenStudy (anonymous):

well i think that the question doesn't make sense, does it?

OpenStudy (anonymous):

because the word "common" means that there must be more than one

OpenStudy (amistre64):

the term common is redundant in it i agree

OpenStudy (anonymous):

no it is not redundant. math eschews verbiage, so no redundant words

OpenStudy (anonymous):

mii choices is a.(x-3)(x+1) b. (x-1)(x-2) c.(x+3)(x-1) d.(x-2)(x+1)

OpenStudy (anonymous):

simple point i am trying to make is that you cannot ask for the greatest common factor of \[40\times 5+14\times 5+5\] it makes no sense

OpenStudy (amistre64):

we get to problems like this: x^2-2x-3 :and the kids dont understand what a factor or a product is. 4 * 3 = 12 4 and 3 are factors of 12 12 is the product of 4*3

OpenStudy (amistre64):

if you have choices; multiply them together to see which one matches

OpenStudy (anonymous):

you cannot ask for the greatest common factor of \[40x+14x+x\] and say that the answer is x

OpenStudy (anonymous):

so question makes no sense at all

OpenStudy (amistre64):

lol ... x may or may not be greatest; but in that case it is common

OpenStudy (anonymous):

it is not! x is the greatest common factor of \[40x,14x,\text{and }x\]

OpenStudy (amistre64):

the clues in your problem suggest that we are subtracting and the bigger number is negative; so id go with a

OpenStudy (anonymous):

any more than 5 is the greatest common factor of 275

OpenStudy (amistre64):

x^2 -2x -3 ^ ^ ^ subtraction ^ bigger factor gets the negative id say "a" is the best bet